### v1

# bibliography.bib

Abstract
Autonomous vehicles (AV) offer a cost-effective solution for scientific missions such as underwater tracking. Recently, reinforcement learning (RL) has emerged as a powerful method for controlling AVs in complex marine environments. However, scaling these techniques to a fleet—essential for multi-target tracking or targets with rapid, unpredictable motion—presents significant computational challenges. Multi-Agent Reinforcement Learning (MARL) is notoriously sample-inefficient, and while high-fidelity simulators like Gazebo’s LRAUV provide 100× faster-than-real-time single-robot simulations, they offer no significant speedup for multi-vehicle scenarios, making MARL training impractical. To address these limitations, we propose an iterative distillation method that transfers high-fidelity simulations into a simplified, GPU-accelerated environment while preserving high-level dynamics. This approach achieves up to a 30,000× speedup over Gazebo through parallelization, enabling efficient training via end-to-end GPU acceleration. Additionally, we introduce a novel Transformer-based architecture (TransfMAPPO) that learns multi-agent policies invariant to the number of agents and targets, significantly improving sample efficiency. Following large-scale curriculum learning conducted entirely on GPU, we perform extensive evaluations in Gazebo, demonstrating that our method maintains tracking errors below 5 meters over extended durations, even in the presence of multiple fast-moving targets. This work bridges the gap between large-scale MARL training and high-fidelity deployment, providing a scalable framework for autonomous fleet control in real-world sea missions.
I INTRODUCTION
Figure 1: Overview of our Training and Evaluation Pipeline.
Figure 2: Five agents trained in the GPU simplified environment follow in Gazebo simulator five fast targets over several kms. Video: https://mttga.github.io/posts/pylrauv/images/5v5.gif
Underwater tracking (UT) is essential to advance marine research missions, such as tracking marine species and oceanographic phenomena [zarokanellos2022frontal, masmitja2020mobile], monitoring large-scale data collection using autonomous underwater vehicles (AUV) [Yanwu2021], and managing marine protected areas [vigo2021spatial]. However, underwater communication systems face challenges such as low reliability and bandwidth, as GPS is ineffective [heidemann2012underwater]. Acoustic tracking using AUVs or autonomous surface vehicles (ASVs) offers a promising solution, improving efficiency and reducing costs compared to traditional fixed equipment [masmitja2020mobile, Yanwu2021]. Yet, challenges like unreliable communication, complex environment dynamics, and energy constraints necessitate advanced motion planning to enhance AVs’ tracking capabilities.
Reinforcement learning (RL) enables AVs to learn optimal navigation planning through trial and error, offering dynamic responses to environmental conditions compared to pre-programmed approaches [cote2019characterizing]. Recent work by [ivanscience] demonstrated the potential of RL by training a Soft Actor-Critic (SAC) agent in a simulator to control an ASV tracking an underwater target using range-only acoustic data. While successful, the model was limited to a single target moving linearly at modest speeds (up to 0.3 times the agent’s velocity).
Scaling this approach to more complex scenarios requires multi-agent systems. Multi-Agent Reinforcement Learning (MARL) could enable learning an optimal cooperative tracking policy by sharing tracking data collected simultaneously from different locations. However, deploying such a complex system in the sea requires preliminary extensive testing. Simulators like Gazebo’s Long-Range AUV simulator (LRAUV Sim) [lrauv_sim_icra], which models hydrodynamics, acoustic communication, and marine sensors, are crucial tools in this phase. Unfortunately, LRAUV Sim offers only a  10x faster-than-real-time speedup in multi-robot simulations, and its C++-based control complicates integration with RL methods, meaning that MARL cannot be applied directly in such a simulator.
To advance multi-agent systems for underwater tracking, we therefore make the following contributions:
  • PyLrauv  [^1] https://github.com/mttga/pylrauv: An open-source Python package using ROS2 [ros2] to control LRAUV Sim with a Gym-like interface. PyLrauv enables seamless control and observation of a variable number of agents and targets while ensuring high-fidelity simulation via Gazebo. It also supports direct deployment in real missions.
  • Environment Distillation: A method to create a simplified, GPU-accelerated environment using JAX, achieving up to 30,000x speedup over Gazebo and allowing to train effective policies in minutes. This distilled environment, which we integrate into the popular MARL library JaxMARL [flair2023jaxmarl]  [^2] https://github.com/FLAIROx/JaxMARL, allows rapid training while maintaining high-level equivalence to PyLrauv.
  • TransfMAPPO: A novel method integrates Transformers into MAPPO [mappo] to train policies independent of the number of entities in the environment. TransfMAPPO enables a curriculum learning (CL) pipeline that learns robust policies in complex multi-agent scenarios, where traditional from-scratch methods like MAPPO are ineffective.
  • High-Fidelity Evaluation: We demonstrate our pipeline effectiveness by training TransfMAPPO purely on GPU on our distilled environment and testing it in high-fidelity simulations using PyLrauv: our trained models can track erratically moving targets at 0.8 times their speed and multiple targets simultaneously. We open-source TransfMAPPO weights for use in LRAUV Sim and real missions via PyLrauv.
II RELATED WORK
Early research in UT with autonomous vehicles has relied on traditional control strategies for AVs [heidemann2012underwater, almeida2017], with more recent work incorporating reinforcement learning (RL) to improve navigation and tracking robustness [ivanscience]. However, these studies have mainly focused on single-agent scenarios or environments with limited target dynamics.
MARL has emerged as a natural extension of RL for learning cooperative policies, yet its sample inefficiency and scalability remain challenging. Transformer-based methods, such as TransfQMix [transfqmix], decouple policy learning from the number of entities by leveraging specialized transformer modules, allowing, in principle, the training of the same network in progressively more challenging multi-agent scenarios. Our TransfMAPPO architecture is inspired by TransfQMix, but uses MAPPO instead of DQN [dqn2016] as the RL backbone due to its natural compatibility with parallelized environments. Moreover, TransfMAPPO does not use a centralized hypernetwork but rather a centralized critic, simplifying the model structure. Finally, TransfMAPPO is also applicable to continuous action spaces, while TransfQMix is limited to discrete.
Frameworks like JaxMARL [flair2023jaxmarl] have demonstrated the benefits of GPU-accelerated MARL, albeit in relatively simple environments, while packages such as IsaacGym [isaacgym] provide more advanced GPU-based simulation of real robots but do not offer the hydrodynamic models and communication protocols necessary to bridge the sim-to-real gap.
Our pipeline can be associated with model-based RL [model_based_1, model_based_2, model_based_3], which aim to learn a world model to improve and speed up RL training. Our method creates a simplified, GPU-accelerated physical model from a high-fidelity simulator, but instead of focusing on step-wise alignment between the world model and the real world, it focuses on preserving high-level dynamics at a different time scale with respect to a more complex simulator. Moreover, our distilled environment is not accessible by our learning algorithm, as is usually the case in model-based RL; instead, we simply use model-free RL on top of it.
III BACKGROUND
III-A Underwater Tracking with Autonomous Vehicles
Underwater acoustic tracking with autonomous vehicles is a member of the family of Range Only Single Beacon (ROSB) localization techniques [rosb, av_range_only], which use range-only acoustic observations gathered over time to locate a target. For static targets, the trilateration problem can be linearised by least squares (LS) algorithms, which assume static conditions and incur errors when targets move [ThesisLisboa]. For dynamic scenarios, particle filters (PF) use Bayesian estimation, representing potential target states with weighted particles updated via motion models and resampled based on measurement likelihoods [ROSB1, ROSB2].
III-B Multi-Agent Reinforcement Learning
Cooperative multi-agent tasks are formalized as decentralized partially observable Markov decision processes (Dec-POMDP) [dec-Pompd], defined by the tuple $G=\langle S,\mathbf{U},P,r,Z,O,n,\gamma\rangle$, where $S$ is the global state space, $\mathbf{U}=U^{n}$ is the joint action space for $n$ agents (each selecting actions $u^{a}\in U$), $P(s^{\prime}\mid s,\mathbf{u})$ is the state transition function, $r(s,\mathbf{u})$ is the shared reward function, $Z$ is the observation space for each agent, $O(s,a)$ defines the observation $z^{a}\in Z$ for agent $a$, and $\gamma\in[0,1)$ is the discount factor. At each timestep, agent $a$ receives a partial observation $z^{a}=O(s,a)$ and maintains an action-observation history $\tau^{a}=(z_{0}^{a},u_{0}^{a},\dots,z_{t}^{a})\in\mathcal{T}$. The agent’s policy $\pi^{a}(u^{a}\mid\tau^{a}):\mathcal{T}\times U\rightarrow[0,1]$ maps this history to a distribution over actions. The agents collectively aim to maximize the expected discounted return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$. A key paradigm in MARL is centralized training with decentralized execution (CTDE) [rashid2018qmixmonotonicvaluefunction], where agents leverage centralized information during training but act independently during execution. This allows the use of a centralized critic that estimates the value function $V(s)$ using the global state $s$, while policies $\pi^{a}$ remain decentralized, depending only on local histories $\tau^{a}$.
MAPPO [mappo] extends Proximal Policy Optimization (PPO) [ppo] to multi-agent settings. Each agent’s policy (actor) $\pi_{\theta}^{a}$ and centralized value function (critic) $V_{\phi}$ are optimized jointly. The actor is updated using the clipped surrogate objective:
$\mathcal{L}^{CLIP}(\theta)=\mathbb{E}_{t}\left[\min\left(\rho_{t}(\theta)\hat{A}_{t},\text{clip}(\rho_{t}(\theta),1-\epsilon,1+\epsilon)\hat{A}_{t}\right)\right],$
where $\rho_{t}(\theta)=\frac{\pi_{\theta}(u_{t}\mid\tau_{t})}{\pi_{\theta_{\text{old}}}(u_{t}\mid\tau_{t})}$ is the probability ratio, $\hat{A}_{t}$ is the advantage estimate, and $\epsilon$ is a clipping hyperparameter. The critic minimizes the temporal difference error:
$\mathcal{L}^{VF}(\phi)=\mathbb{E}_{t}\left[\left(V_{\phi}(s_{t})-\hat{R}_{t}\right)^{2}\right],$
where $\hat{R}_{t}$ is the discounted return. During execution, only the actors are used, enabling decentralized control.
IV METHOD
IV-A PyLrauv
PyLrauv provides a Python interface to LRAUV Sim by leveraging ROS 2 [ros2]. This is achieved by integrating a set of additional communication handlers into the official Gazebo C++ library, ros_gz [^3] https://github.com/gazebosim/ros_gz. The integration allows control of the LRAUV robots with python using rclpy  [^4] https://github.com/ros2/rclpy. We have developed a set of Python controllers to send commands to the LRAUV vehicles in Gazebo and observe their published states. Since the actual LRAUV vehicles use the same technology, the controllers provided within PyLrauv can also be used to control real LRAUV robots. Controlling multiple vehicles simulated in Gazebo in a Gym-like fashion is straightforward, as demonstrated by the following code:
1from pylrauv import UTrackingEnv
2
3env = UTrackingEnv(num_agents=2, num_targets=2)
4obs, state = env.reset()
5actions = {’agent_1’: 0, ’agent_2’: 0}
6obs, state, reward, done = env.step(actions, step_time=30)
IV-B UTracking Environment
UTrackingEnv is a Gym-like environment managing agent and target commands, communication, tracking, and RL tasks like observation building and reward calculations. Agents communicate via the Gazebo LRAUV Acoustic Comms Plugin and collect target ranges using the LRAUV Range Bearing Plugin. At the start of an episode, agents are spawned on the sea surface and targets at random depths, ensuring a minimum distance between each entity of 50 meters and maximum of 200. Both agents and targets are LRAUV vehicles. Each step lasts 30 seconds in our experiments, at the beginning of which agents listen for range signals from targets, and at the end broadcast their locations and observations to other agents. Each agent maintain its own tracking model (LS or PF) for each target, updated with all the information received. This ensures a genuine decentralization of the system. Targets are tracked in 2D space, as depth is typically known [ivanscience].
Observation Space
The agents receive their partial observation of the environment at the end of the step phase, which includes the relative distances to all other entities in the 3D space. For other agents, these distances are calculated using the information they have received from them (if available), and for targets, the distances are calculated from the tracking predictions generated by their own tracking modules. [ivanscience] incorporates the range observations into observation vectors, which may be helpful, but may also cause agents to overfit to training dynamics and overlook tracking information. To prevent this, we limit observations to tracking-derived values, ensuring agents rely on tracking data and its inherent errors. The environment also returns the global state of the environment, which includes the true 3D position of each entity, its velocity, and its direction.
Reward Function
[…]
$\small r_{i}^{tracking}=\begin{cases}1,&e_{i}<\epsilon_{\text{min}},\\ \exp\left(-\dfrac{2t}{1-t}\right),&\epsilon_{\text{min}}\leq e_{i}\leq\epsilon_{\text{max}},\\ 0,&e_{i}>\epsilon_{\text{max}},\end{cases}$ (1)
with $t=(e_{i}-\epsilon_{\text{min}})/({\epsilon_{\text{max}}-\epsilon_{\text{min}}})$. We set $\epsilon_{\text{min}}=10\,\text{m}$ and $\epsilon_{\text{max}}=50\,\text{m}$. The exponential decay encourages precise tracking, distinguishing more clearly states where the tracking was closer to the ideal tracking error. The follow reward is based on the proper distribution of agents with respect to the targets. This reward encourages a perfect distribution across targets for the agents, which is known to be the best policy when $N$ agents need to follow (rapid enough) $N$ targets. For target $i$, if the minimum distance between it and any agent, $d_{i}$, is less than specified threshold, $d_{\min}$, then the target is considered successfully followed. The reward is computed as
$r_{i}^{follow}=\mathbb{I}\left\{d_{i}\leq\,d_{\min}\right\},$ (2)
[…]
$r=\frac{1}{N}\sum_{i=1}^{N}r_{i},$ (3)
where $N$ is the number of targets to track, i.e. the reward is normalized by the number of targets to be always in the interval $[0,1]$, so that learned value functions can be transferred across scenarios with different numbers of targets. In addition, a crash penalty is always applied if the distance between any two agents falls below a minimum valid distance, $d_{safe}$, i.e., $r=-1\quad\text{if}\quad\min_{i\neq j}\{d_{ij}\}<d_{safe}.$
Actions Space
To simplify the action space of our multi-agent system, we keep the agents’ velocity constant and allow them to control only the rudder. The rudder is discretized into five values corresponding to equidistant angles ranging from $-0.24$ to $0.24$ radians. Agents can adjust the rudder only to the next closest two angles, preventing abrupt changes in direction.
Termination
The episode ends when the maximum number of steps is reached. The environment is not truncated when crashes occur or targets are lost, as we experimentally observed that truncating episodes would create training instabilities.
Why UTracking is a difficult environment?
  • Partial observability: Agents detect targets only within 450 m via noisy acoustic signals and rely on intermittent and unstable communication (1500 m range) for shared tracking data.
  • Stochasticity: Sensor noise, communication dropouts, environmental perturbations, and unpredictable target motions amplify uncertainty.
  • Coordination challenge: Agents must coordinate almost perfectly especially when tracking multiple rapid-moving targets, since a lack of coordination could result in rapidly loosing a target.
  • Long horizons: Hour-scale missions (corresponding to thousands of steps) require learning delayed consequences of actions, which can be difficult in Multi-Agent scenarios given their non-stationarity. Moreover, they need to be robust to error accumulation in tracking models.
IV-C Environment Distillation: JaxLrauv UTracking
LRAUV Sim simulates the dynamics of a LRAUV robot at the millisecond level. However, LRAUV vehicles take actions only every few seconds to minutes. Morover, the tracking missions we are interested in span several hours. For this reason, it would be beneficial to simulate the environment state at a much larger time scale. Formally, we are interested in modeling the new position in the 2D plane $\mathbf{p}_{t+1}\in\mathbb{R}^{2}$ after a delta time $\delta_{t}$ given the actual robot position $\mathbf{p}_{t}$, absolute speed $v$ and rudder angle $\gamma$. Instead of reinventing the wheel and re-building an accurate physical model which is already provided by Gazebo, we want to approximate the change in the robots’ heading $\delta_{\psi}$ in order to use a simplified trajectory model, i.e. $\mathbf{p}_{t+1}=\mathbf{p}_{t}+v\delta_{t}\begin{bmatrix}\cos(\psi_{t}+\delta_{\psi}),\sin(\psi_{t}+\delta_{\psi})\end{bmatrix}$. Fixing $v$ and $\delta_{t}$, we parametrize $\delta_{\gamma}$ in respect to the rudder angle $\theta(\gamma)$. To approximate $\theta(\gamma)$ we follow an iterative distillation process: we collect trajectories in PyLrauv, build a simplified model, train an agent on that, and collect new data using the trained agent to improve the model distillation. We found that using an ensemble of linear models it is possible to approximate $\theta(\gamma)$ with a mean absolute error below 0.015 radians and a global $R^{2}$ score of 0.99. Such a simple model fails to grasp long time-dependency, and diverges when used autoregressively to generate long trajectories. Nevertheless, it performs well for shorter trajectories and is extremely faster than, for example, a $LSTM$. To improve the robustness of the system, we introduce Gaussian noise with a standard deviation of 0.02 radians, forcing the trained agent to handle trajectory uncertainty. In addition, we equip the agents with recurrent mechanisms. We found that training in this way in the simplified environment, the learned policies are directly transferable to Gazebo.
The rest of the environment, including measurement noise, the probability of communication loss, partial observability and collision handling, is implemented manually following the observed statistics from the collected data. The final simplified environment is implemented in JAX to enable an end-to-end GPU-based RL pipeline, and is called JaxLrauv UTracking. The speedups achieved in different configurations of agents and targets are presented in LABEL:tab:speedup. These speedups allow to train a single-target tracking model that can be directly deployed in LRAUV Sim in as little as 10 minutes.
TABLE I: Speedups Obtained with the Distilled Environment.
Config Pylrauv JaxLrauv 1 Envs JaxLrauv 128 Envs JaxLrauv 1024 Envs
SPS SPS Speedup SPS Speedup SPS Speedup
1A,1T 2.7 1289 477 68508 25352 81686 30229
2A,2T 1.0 1068 1084 19712 20017 21534 21867
3A,3T 0.4 1053 2439 9128 21139 9786 22663
4A,4T 0.3 1020 3571 5180 18125 5450 19070
5A,5T 0.2 936 4751 3469 17607 3574 18142
IV-D Vectorized Particle Filter
While PF represent a state-of-the-art method for underwater tracking, their computational expense renders them inefficient for training, particularly in multi-agent scenarios. To mitigate this, [ivanscience] suggests employing Least Squares (LS) during training and reserving PF for testing. However, LS is only stable for static or slow-moving targets, limiting its scalability for faster targets and potentially intensifying the inherent non-stationarity of MARL. To facilitate efficient training, we implement a PF in JAX, utilizing vmap for GPU parallelization. By nesting vmapped functions, during training we update all particles across targets, agents, and parallel environments in a single pass, reducing the speed loss to only 30% compared to LS. As demonstrated in Figure 3, training with LS fails for targets moving faster than one-third of the agent’s velocity, whereas PF maintains stability.
Figure 3: Using Particle Filter or Least Squares to train an agent to follow a target at different velocities.
IV-E TransfMAPPO
The complexity of training multi-agent system with RL typically scales with the number of agents involved. To address this, we introduce TransfMAPPO, inspired by TransfQMix [transfqmix], which maintains the core concepts of the latter but uses PPO as the RL backbone instead of DQN, since PPO is more natural for vectorized environments. In TransfMAPPO, we implement a Transformer Actor and Critic that process the features describing the observed entities as vertices of a coordination latent graph learned through self-attention. The resulting policy and Value function are therefore invariant from the number of entities observed, allowing curriculum learning procedure described in subsection IV-F. More formally, the Transformer Critic observes the state of the environment represented as $n$ entities each one described by $z$ features, which in our case are the position, velocity and angles of the agents and targets. The Transformer Agent also observes the $n$ entities, but from its own partial observability perspective, which is constraint by the acoustic communication protocols defined in section IV. An overview of TransfMAPPO can be seen in Figure 4. Following [gallici2024simplifyingdeeptemporaldifference], we found that normalizing the entire network with LayerNormalization greatly improves stability and performances.
Figure 4: Overview of TransfMAPPO architecture.
IV-F Curriculum Learning
Taking advantage of our Transformer-based architecture, we employ curriculum learning to progressively learn Multi-Agent general policies starting from a Single-Agent policy (see Figure 5). We begin with a large pre-training phase in which we train a single agent to maximize the tracking reward described in subsection IV-A with respect to an erratic-moving fast target (0.6x of its speed). We train for $10^{9}$ timesteps, corresponding to 1.5 days of compute (that would take months in Gazebo). In this phase, we maintain a short horizon (128 steps) before resetting the environment, as it significantly aids learning. The resulting policy is able to track a fast target with an error below 35 meters, but it fails to generalize to longer horizons. We therefore fine-tune the policy in progressively longer episodes (256, 512, and 1024 steps) for $10^{7}$ timesteps, until we observe that the policy can perfectly track a fast target for more than 10k steps, corresponding to 3.4 days of real-time tracking.
From this horizon-invariant base model, we transition to a multi-agent phase, where we fine-tune the policy for cooperative tasks. In particular, we consider two scenarios: (1) $n$ agents tracking $n$ targets, where we use the follow reward, and (2) $n$ agents tracking one very fast target (0.8x of their velocity), maximizing the same tracking reward used in the single-agent phase. Given that these two scenarios pose different cooperation and credit assignment problems, we reset the Critic parameters and maintain separate sets of fine-tuned agents for the two multi-agent fine-tuning branches, resulting in a final multi-target follow model and a multi-robot tracking model. In both cases, we fine-tune the model to cooperate with progressively more agents (5 and 3 at maximum, respectively) for a maximum $20^{8}$ timesteps. We observe that both resulting models can cooperate with a variable number of agents. In this phase, we also note that maintaining long environment horizons negatively affects training, so we reduce the horizon back to 256 steps. Unfortunately, while the resulting multi-agent tracking model retains the horizon invariance of the pre-trained single-agent model, the multi-target follow model loses some invariance in comparison.
Figure 5: Overview of our Curriculum Learning Procedure.
V EXPERIMENTS
All the training was performed on a single H100 GPU of MareNostrum 5  [^5] https://www.bsc.es/ca/marenostrum/marenostrum-5, while the Gazebo simulator experiments run on a Intel Sapphire Rapids 8460Y CPU. All the episode returns are normalized for the horizon length to be presented in the interval $[0,1]$. All the training results are averaged across 5 seeds, except for the curriculum models. All the figures representing agents trajectories are produced with data collected in the Gazebo simulator.
V-A Multi-Robot Tracking
The effectiveness of our curriculum method is demonstrated in Figure 6, where we compare the training curve of fine-tuning the model for the task of tracking a very fast target (0.8x the agents’ velocity) with 3 agents against training from scratch on the same task. The superior performance of curriculum learning is also evident in the reduced tracking error. In the same plot, we also include the training curve of single-agent PPO, which achieves less than half the returns of the multi-agent system.
Figure 6: Training multiple agents to track a very fast target.
Figure 7: Multi-Robot Tracking Evaluation. The highglighted area represents the values seen in training.
To validate our final multi-robot tracking model based on TransfMAPPO, we conduct a comprehensive validation analysis using JaxLrauv Figure 10. Every data point in the validation phase is averaged across 1000 episodes, each with a length of 256 steps, following a target moving at 0.5x the agents’ velocity (unless otherwise specified). Specifically, we test how the tracking error is affected under three conditions: (1) when the target moves faster, (2) when its motion becomes more unpredictable (defined by the frequency of direction changes), and (3) when the agents’ motion is perturbed (simulating external forces such as ocean currents). In the last sub-figure we illustrate how the multi-agent system’s performance scales naturally with the number of agents used to track the target. Notably, even using only 2 agents significantly enhances the system’s robustness compared to a single-agent approach. We also highlight that the tracking error decreases monotonically with the number of agents, with 3 agents representing the sweet spot, achieving negligible collision probability while maintaining optimal tracking performance.
V-B Multi-Target Tracking
Similarly, our curriculum method achieves remarkable advantages compared to training MAPPO from scratch in the multi-target tracking scenario. In Figure 8, we present the training curves of training MAPPO from scratch to track moderately fast targets (up to 0.5x the agents’ velocity). MAPPO can learn to follow barely half of the targets only in the 2-agent-2-target scenario, while our curriculum approach successfully follows more than 90% of the targets on average, even in the 5-agent-5-target scenario.
Figure 8: Training Multiple Agents to Follow Fast Targets.
To further test the effectiveness of TransfMAPPO, we train it from scratch alongside MAPPO in a simpler configuration where targets move slower (0.3x the agents’s speed) and in a more predictable manner. In this case, TransfMAPPO and MAPPO achieve similar scores with 2 agents, but MAPPO’s performance drops significantly with 3 agents. Surprisingly, TransfMAPPO learns very effectively even from scratch in the 5-agent-5-target scenario, confirming the sample efficiency obtained by decoupling a cooperation policy using Transformers.
Figure 9: Training Multiple Agents to Follow Slower Targets.
V-C Results in Gazebo
We conducted a series of final experiments in Gazebo to evaluate the performance of our trained agents in a high-fidelity simulation environment. We collected data over 50 episodes using the models obtained at the conclusion of the curriculum training. Specifically, we performed the following tests:
  1. 1.
    Tracking a slow target (0.3$\times$ the agent’s velocity) for 300 steps using a single robot trained for 10 minutes in JaxLrauv.
  2. 2.
    Tracking a fast target (0.6$\times$ the agent’s velocity) for 1000 steps (equivalent to over 8 hours of real-time tracking) using our horizon-invariant base model.
  3. 3.
    Tracking a very fast target (0.8$\times$ the agent’s velocity) for 1000 steps using a multi-robot tracking fine-tuned model with 3 agents.
  4. 4.
    Tracking 3 targets simultaneously moving at 0.5$\times$ the agents’ velocity using a multi-target model for 300 steps.
  5. 5.
    Tracking 5 targets simultaneously moving at 0.5$\times$ the agents’ velocity using a multi-target model for 300 steps.
For the multi-target experiments, we reduced the number of steps to 300 due to the computational burden of simulating multiple robots in Gazebo. The results of these experiments are presented in Table II. Our findings demonstrate that even with only 10 minutes of training, we were able to develop an effective model capable of maintaining target tracking without losing the target. The multi-robot tracking model achieved a promising average tracking error of 3 meters, even when following a very fast target. This performance highlights a clear advantage over single-agent trackers following slower targets. Notably, the average agent-target distance increased when using multiple robots, reflecting a broader distribution of robots around the target space. The multi-target models performed as expected, with a small 5% probability of losing a single landmark. However, the multi-agent models exhibited an increased probability of collisions. This issue is likely attributable to the fact that the models were trained to avoid collisions only every 30 seconds, which can result in intra-step collisions in Gazebo. This underscores a critical area for future work to enhance the systems’ safety before deploying them in real-world marine missions.
TABLE II: Experimental Results in Gazebo
Config Avg. Agent-Target Distance Av. Tracking Error Prob. Collision (%) Prob. Lost (%)
1A,1T (Slow, 10Mins Train) 59.10 ± 32.29 5.89 ± 3.7 0.00 0.00
1A,1T (Fast) 121.01 ± 62.11 20.33 ± 28.38 0.00 7.14
3A,1T (Very Fast) 217.64 ± 89.85 3.03 ± 6.17 15.32 0.00
3A,3T (Moderate) 45.30 ± 26.11 5.29 ± 4.88 5.00 5.00
5A,5T (Moderate) 49.24 ± 30.92 4.25 ± 4.57 10.00 5.26
VI Conclusions
We presented a method for scaling up Multi-Agent Reinforcement Learning (MARL) in the context of underwater tracking, demonstrating how it is possible to efficiently train MARL models purely on GPUs and deploy them in high-fidelity simulators, where MARL training would otherwise be extremely expensive or infeasible. Although this method relied on manual adjustments during the environment distillation phase, our overall pipeline represents a promising direction for closing the sim-to-real gap in MARL. A natural next step for this work is to enhance the safety of the multi-agent system and test it in real-world sea missions.
Figure 10: Multi-agent episodes collected in Gazebo. (a): Single-agent tracking, target velocity 0.66× of agent velocity. Video. (b): Multi-agent tracking, target velocity 0.88x of agents velocity. Video.
Figure 11: Agents trajectory according to target speed. Notice the change in the curvature of the agent’s trajectory to optimally track targets moving at different speeds.
Figure 12: Coordination when tracking a very fast target. Notice how agents react to a rapid change of direction of the target by waiting each other in order to not loose communication and then rapidly moving together.
Figure 13: Coordination in multi-target setting. Notice how agents resolve the ”traffic” problem by making circles to wait for other agents’ passing, and one target is dynamically reassigned from agent 3 to agent 1.
ACKNOWLEDGMENT
This work acknowledges the Spanish Ministerio de Ciencia, Innovación y Universidades (BITER-ECO: PID2020-114732RBC31). This work is part of DIGI4ECO, which has received funding from the European Union’s Horizon Europe programme under grant agreement No 101112883. I. M. received financial support from the MCIN/AEI/10.13039/501100011033 and FSE+, grant agreement No RYC2022-038056-I. M. G. was partially founded by the FPI-UPC Santander Scholarship FPI-UPC_93. This work also acknowledges the ’Severo Ochoa Centre of Excellence’ accreditation (CEX2019-000928-S) and the Barcelona Super Computer Center.
[…]

### v2

@@
-bibliography.bib
+Scaling Multi-Agent Reinforcement Learning for Underwater Acoustic Tracking via Autonomous Vehicles
 Abstract
-Autonomous vehicles (AV) offer a cost-effective solution for scientific missions such as underwater tracking.
-Recently, reinforcement learning (RL) has emerged as a powerful method for controlling AVs in complex marine environments.
-However, scaling these techniques to a fleet—essential for multi-target tracking or targets with rapid, unpredictable motion—presents significant computational challenges.
-Multi-Agent Reinforcement Learning (MARL) is notoriously sample-inefficient, and while high-fidelity simulators like Gazebo’s LRAUV provide 100× faster-than-real-time single-robot simulations, they offer no significant speedup for multi-vehicle scenarios, making MARL training impractical.
-To address these limitations, we propose an iterative distillation method that transfers high-fidelity simulations into a simplified, GPU-accelerated environment while preserving high-level dynamics.
-This approach achieves up to a 30,000× speedup over Gazebo through parallelization, enabling efficient training via end-to-end GPU acceleration.
-Additionally, we introduce a novel Transformer-based architecture (TransfMAPPO) that learns multi-agent policies invariant to the number of agents and targets, significantly improving sample efficiency.
-Following large-scale curriculum learning conducted entirely on GPU, we perform extensive evaluations in Gazebo, demonstrating that our method maintains tracking errors below 5 meters over extended durations, even in the presence of multiple fast-moving targets.
-This work bridges the gap between large-scale MARL training and high-fidelity deployment, providing a scalable framework for autonomous fleet control in real-world sea missions.
+Autonomous vehicles (AVs) offer a cost-effective solution for scientific missions such as underwater tracking.
+Reinforcement learning (RL) has emerged as a powerful method for controlling AVs, but scaling to fleets (essential for multi-target tracking or rapidly moving targets) is challenging.
+Multi-Agent RL (MARL) is notoriously sample-inefficient, and while high-fidelity simulators like Gazebo’s LRAUV provide up to 100× faster-than-real-time single-robot simulations, they offer little speedup in multi-vehicle scenarios, making MARL training impractical.
+Yet, high-fidelity simulation is crucial to test complex policies and close the sim-to-real gap.
+To address these limitations, we develop a GPU-accelerated environment that achieves up to 30,000× speedup over Gazebo while preserving its dynamics.
+This enables fast, end-to-end GPU training and seamless transfer to Gazebo for evaluation.
+We also introduce a Transformer-based architecture (TransfMAPPO) that learns policies invariant to fleet size and number of targets, enabling curriculum learning to train larger fleets on increasingly complex scenarios.
+After large-scale GPU training, we perform extensive evaluations in Gazebo, showing our method maintains tracking errors below 5m even with multiple fast-moving targets.
 I INTRODUCTION
+Underwater tracking (UT) is essential to advance marine research missions, such as tracking marine species and oceanographic phenomena [1, 2], monitoring large-scale data collection using autonomous underwater vehicles (AUV) [3], and managing marine protected areas [4].
+However, underwater communication systems face challenges such as low reliability and bandwidth, as GPS is ineffective [5].
+Acoustic tracking using AUVs or autonomous surface vehicles (ASVs) offers a promising solution, improving efficiency and reducing costs compared to traditional fixed equipment [2, 3].
+Yet, challenges like unreliable communication, complex environment dynamics, and energy constraints necessitate advanced motion planning to enhance AVs’ tracking capabilities.
 Figure 1: Overview of our training and evaluation pipeline.
+PyLrauv $(a)$ is a new Python package to control multiple robots in the C++ high-fidelity LRAUV simulator.
+JaxLrauv $(b)$ is a GPU-accelerated, simplified environment that supports massive parallelization while preserving the dynamics of the LRAUV simulator.
+TransfMAPPO $(c)$ employs transformers to train progressively larger fleets of vehicles to coordinate via curriculum learning.
+The final policies trained in JaxLrauv with TransfMAPPO are then evaluated in the realistic LRAUV simulator prior to real-world deployment.
 Figure 2: Five agents trained in the GPU simplified environment follow in Gazebo simulator five fast targets over several kms. Video: https://mttga.github.io/posts/pylrauv/images/5v5.gif
-Underwater tracking (UT) is essential to advance marine research missions, such as tracking marine species and oceanographic phenomena [zarokanellos2022frontal, masmitja2020mobile], monitoring large-scale data collection using autonomous underwater vehicles (AUV) [Yanwu2021], and managing marine protected areas [vigo2021spatial].
-However, underwater communication systems face challenges such as low reliability and bandwidth, as GPS is ineffective [heidemann2012underwater].
-Acoustic tracking using AUVs or autonomous surface vehicles (ASVs) offers a promising solution, improving efficiency and reducing costs compared to traditional fixed equipment [masmitja2020mobile, Yanwu2021].
-Yet, challenges like unreliable communication, complex environment dynamics, and energy constraints necessitate advanced motion planning to enhance AVs’ tracking capabilities.
-Reinforcement learning (RL) enables AVs to learn optimal navigation planning through trial and error, offering dynamic responses to environmental conditions compared to pre-programmed approaches [cote2019characterizing].
-Recent work by [ivanscience] demonstrated the potential of RL by training a Soft Actor-Critic (SAC) agent in a simulator to control an ASV tracking an underwater target using range-only acoustic data.
-While successful, the model was limited to a single target moving linearly at modest speeds (up to 0.3 times the agent’s velocity).
-Scaling this approach to more complex scenarios requires multi-agent systems.
-Multi-Agent Reinforcement Learning (MARL) could enable learning an optimal cooperative tracking policy by sharing tracking data collected simultaneously from different locations.
-However, deploying such a complex system in the sea requires preliminary extensive testing.
-Simulators like Gazebo’s Long-Range AUV simulator (LRAUV Sim) [lrauv_sim_icra], which models hydrodynamics, acoustic communication, and marine sensors, are crucial tools in this phase.
-Unfortunately, LRAUV Sim offers only a  10x faster-than-real-time speedup in multi-robot simulations, and its C++-based control complicates integration with RL methods, meaning that MARL cannot be applied directly in such a simulator.
+Reinforcement learning (RL) enables AVs to learn optimal navigation strategies through trial and error, providing dynamic responses to environmental conditions compared to pre-programmed approaches [6].
+Preliminary work by [7] demonstrated the potential of using RL to train an ASV policy for tracking an underwater target using range-only acoustic data.
+More recent studies [8, 9] have shown that Multi-Agent Reinforcement Learning (MARL) can enable cooperative tracking by sharing data collected simultaneously from different locations.
+However, these MARL studies were conducted in very simple, abstract environments.
+Deploying such a complex system in the sea requires extensive preliminary testing.
+Simulators such as Gazebo’s Long-Range AUV simulator (LRAUV Sim) [10], which models hydrodynamics, acoustic communication, and marine sensors of real vehicles, are crucial tools for this phase.
+Unfortunately, LRAUV Sim provides only about a $10\times$ faster-than-real-time speedup in multi-robot simulations, and its C++-based control complicates integration with RL methods, making direct application of MARL impractical in this simulator.
 To advance multi-agent systems for underwater tracking, we therefore make the following contributions:
-  • PyLrauv  [^1] https://github.com/mttga/pylrauv: An open-source Python package using ROS2 [ros2] to control LRAUV Sim with a Gym-like interface.
-PyLrauv enables seamless control and observation of a variable number of agents and targets while ensuring high-fidelity simulation via Gazebo.
-It also supports direct deployment in real missions.
-  • Environment Distillation: A method to create a simplified, GPU-accelerated environment using JAX, achieving up to 30,000x speedup over Gazebo and allowing to train effective policies in minutes.
-This distilled environment, which we integrate into the popular MARL library JaxMARL [flair2023jaxmarl]  [^2] https://github.com/FLAIROx/JaxMARL, allows rapid training while maintaining high-level equivalence to PyLrauv.
-  • TransfMAPPO: A novel method integrates Transformers into MAPPO [mappo] to train policies independent of the number of entities in the environment.
-TransfMAPPO enables a curriculum learning (CL) pipeline that learns robust policies in complex multi-agent scenarios, where traditional from-scratch methods like MAPPO are ineffective.
-  • High-Fidelity Evaluation: We demonstrate our pipeline effectiveness by training TransfMAPPO purely on GPU on our distilled environment and testing it in high-fidelity simulations using PyLrauv: our trained models can track erratically moving targets at 0.8 times their speed and multiple targets simultaneously.
-We open-source TransfMAPPO weights for use in LRAUV Sim and real missions via PyLrauv.
+[…]
+  • Agent-Target Invariant Policies with Transformers and Curriculum Learning: By integrating Transformers into a novel variant of MAPPO [13] (TransfMAPPO), we train policies that are invariant to the number of agents and targets.
+This enables a curriculum learning (CL) pipeline that progressively trains larger fleets in increasingly complex multi-target scenarios.
+TransfMAPPO learns robust cooperative policies where traditional from-scratch approaches such as MAPPO fail.
+Thanks to the combination of fast training and curriculum learning, we obtain a general policy capable of tracking up to 5 simultaneous targets with only 5 vehicles (f.i.
+Figure 2), whereas state-of-the-art require up to 12 vehicles to track 4 targets.
 II RELATED WORK
-Early research in UT with autonomous vehicles has relied on traditional control strategies for AVs [heidemann2012underwater, almeida2017], with more recent work incorporating reinforcement learning (RL) to improve navigation and tracking robustness [ivanscience].
-However, these studies have mainly focused on single-agent scenarios or environments with limited target dynamics.
-MARL has emerged as a natural extension of RL for learning cooperative policies, yet its sample inefficiency and scalability remain challenging.
-Transformer-based methods, such as TransfQMix [transfqmix], decouple policy learning from the number of entities by leveraging specialized transformer modules, allowing, in principle, the training of the same network in progressively more challenging multi-agent scenarios.
-Our TransfMAPPO architecture is inspired by TransfQMix, but uses MAPPO instead of DQN [dqn2016] as the RL backbone due to its natural compatibility with parallelized environments.
-Moreover, TransfMAPPO does not use a centralized hypernetwork but rather a centralized critic, simplifying the model structure.
-Finally, TransfMAPPO is also applicable to continuous action spaces, while TransfQMix is limited to discrete.
-Frameworks like JaxMARL [flair2023jaxmarl] have demonstrated the benefits of GPU-accelerated MARL, albeit in relatively simple environments, while packages such as IsaacGym [isaacgym] provide more advanced GPU-based simulation of real robots but do not offer the hydrodynamic models and communication protocols necessary to bridge the sim-to-real gap.
-Our pipeline can be associated with model-based RL [model_based_1, model_based_2, model_based_3], which aim to learn a world model to improve and speed up RL training.
-Our method creates a simplified, GPU-accelerated physical model from a high-fidelity simulator, but instead of focusing on step-wise alignment between the world model and the real world, it focuses on preserving high-level dynamics at a different time scale with respect to a more complex simulator.
-Moreover, our distilled environment is not accessible by our learning algorithm, as is usually the case in model-based RL; instead, we simply use model-free RL on top of it.
+Early research in UT with autonomous vehicles has relied on traditional control strategies for AVs [5, 14].
+A promising new direction was opened by [7], who incorporated RL to improve navigation and tracking robustness, though their work was limited to single vehicles tracking fixed or slowly moving targets.
+More recent studies [9, 8] have explored the use MARL to train cooperative policies for tracking moving targets, but these methods were implemented in highly abstract, simplified environments.
+Our proposed pipeline, in contrast, ensures that learned policies are robust to realistic scenarios.
+Moreover  [9] only demonstrated a cooperative policy with two vehicles tracking a single target.
+Similarly to us, [8] employed an attention mechanism, but did not train team-size invariant policies.
+Instead, they trained fixed configurations such as 4 agents tracking 2 targets, 6 agents and 3 targets, or 12 agents and 4 targets.
+By comparison, our method is able to train policies that generalize to team sizes and target numbers, and we demonstrate robust tracking of 5 targets using only 5 vehicles.
+Neither of these works provide wall-clock training times or official code-bases to obtain them, but they certainly lack environment vectorization and pure-GPU training, allowing us to safely assert that in a comparison our pipeline would be significantly faster and more robust.
+Our TransfMAPPO architecture is inspired by TransfQMix [15], but adopts PPO instead of Q-Learning as the RL backbone due to its natural compatibility with parallelized environments.
+Moreover, TransfMAPPO replaces the centralized hypernetwork with a centralized critic, simplifying the model structure.
+Unlike TransfQMix, TransfMAPPO also supports continuous action spaces, rather than being restricted to discrete ones.
+Several other methods have explored scaling coordination in MARL using attention mechanisms, Graph Neural Networks (GNNs), or Transformers [16, 17, 18, 19], but a detailed comparison with these approaches is beyond the scope of this work.
+The most similar work to ours that combines transformers with MAPPO is Multi-Agent Transformer [20].
+However, in that work transformers are used to model the multi-agent problem as a sequential decision process, whereas in our case we employ them to improve simultaneous coordination.
+Related to our approach, [21, 22] employ curriculum learning to progressively train policies that adapt to larger team sizes, but these methods have not been evaluated in complex robotic scenarios.
 III BACKGROUND
-III-A Underwater Tracking with Autonomous Vehicles
-Underwater acoustic tracking with autonomous vehicles is a member of the family of Range Only Single Beacon (ROSB) localization techniques [rosb, av_range_only], which use range-only acoustic observations gathered over time to locate a target.
-For static targets, the trilateration problem can be linearised by least squares (LS) algorithms, which assume static conditions and incur errors when targets move [ThesisLisboa].
-For dynamic scenarios, particle filters (PF) use Bayesian estimation, representing potential target states with weighted particles updated via motion models and resampled based on measurement likelihoods [ROSB1, ROSB2].
+III-A Underwater Localization with Autonomous Vehicles
+Underwater acoustic tracking with autonomous vehicles is a member of the family of Range Only Single Beacon (ROSB) localization techniques [23, 24], which use range-only acoustic observations gathered over time to locate a target.
+For static targets, the trilateration problem can be linearised by least squares (LS) algorithms, which assume static conditions and incur errors when targets move [25].
+For dynamic scenarios, particle filters (PF) use Bayesian estimation, representing potential target states with weighted particles updated via motion models and resampled based on measurement likelihoods [26, 27].
 III-B Multi-Agent Reinforcement Learning
-Cooperative multi-agent tasks are formalized as decentralized partially observable Markov decision processes (Dec-POMDP) [dec-Pompd], defined by the tuple $G=\langle S,\mathbf{U},P,r,Z,O,n,\gamma\rangle$, where $S$ is the global state space, $\mathbf{U}=U^{n}$ is the joint action space for $n$ agents (each selecting actions $u^{a}\in U$), $P(s^{\prime}\mid s,\mathbf{u})$ is the state transition function, $r(s,\mathbf{u})$ is the shared reward function, $Z$ is the observation space for each agent, $O(s,a)$ defines the observation $z^{a}\in Z$ for agent $a$, and $\gamma\in[0,1)$ is the discount factor.
+Cooperative multi-agent tasks are formalized as decentralized partially observable Markov decision processes (Dec-POMDP) [28], defined by the tuple $G=\langle S,\mathbf{U},P,r,Z,O,n,\gamma\rangle$, where $S$ is the global state space, $\mathbf{U}=U^{n}$ is the joint action space for $n$ agents (each selecting actions $u^{a}\in U$), $P(s^{\prime}\mid s,\mathbf{u})$ is the state transition function, $r(s,\mathbf{u})$ is the shared reward function, $Z$ is the observation space for each agent, $O(s,a)$ defines the observation $z^{a}\in Z$ for agent $a$, and $\gamma\in[0,1)$ is the discount factor.
 At each timestep, agent $a$ receives a partial observation $z^{a}=O(s,a)$ and maintains an action-observation history $\tau^{a}=(z_{0}^{a},u_{0}^{a},\dots,z_{t}^{a})\in\mathcal{T}$.
@@
 The agents collectively aim to maximize the expected discounted return $\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}r_{t}]$.
-A key paradigm in MARL is centralized training with decentralized execution (CTDE) [rashid2018qmixmonotonicvaluefunction], where agents leverage centralized information during training but act independently during execution.
+A key paradigm in MARL is centralized training with decentralized execution (CTDE) [29], where agents leverage centralized information during training but act independently during execution.
 This allows the use of a centralized critic that estimates the value function $V(S)$ using the global state $S$, while policies $\pi^{a}$ remain decentralized, depending only on local histories $\tau^{a}$.
-MAPPO [mappo] extends Proximal Policy Optimization (PPO) [ppo] to multi-agent settings.
+MAPPO [13] extends Proximal Policy Optimization (PPO) [30] to multi-agent settings.
 Each agent’s policy (actor) $\pi_{\theta}^{a}$ and centralized value function (critic) $V_{\phi}$ are optimized jointly.
@@
 where $\rho_{t}(\theta)=\frac{\pi_{\theta}(u_{t}\mid\tau_{t})}{\pi_{\theta_{\text{old}}}(u_{t}\mid\tau_{t})}$ is the probability ratio, $\hat{A}_{t}$ is the advantage estimate, and $\epsilon$ is a clipping hyperparameter.
-The critic minimizes the temporal difference error:
-$\mathcal{L}^{VF}(\phi)=\mathbb{E}_{t}\left[\left(V_{\phi}(s_{t})-\hat{R}_{t}\right)^{2}\right],$
-where $\hat{R}_{t}$ is the discounted return. During execution, only the actors are used, enabling decentralized control.
+The critic minimizes the temporal difference error: $\mathcal{L}^{VF}(\phi)=\mathbb{E}_{t}\left[\left(V_{\phi}(s_{t})-\hat{R}_{t}\right)^{2}\right]$, where $\hat{R}_{t}$ is the discounted $\lambda$ return.
+During execution, only the actors are used, enabling decentralized control.
 IV METHOD
 IV-A PyLrauv
-PyLrauv provides a Python interface to LRAUV Sim by leveraging ROS 2 [ros2].
+PyLrauv provides a Python interface to LRAUV Sim by leveraging ROS 2 [11].
 This is achieved by integrating a set of additional communication handlers into the official Gazebo C++ library, ros_gz [^1] https://github.com/gazebosim/ros_gz.
@@
 6obs, state, reward, done = env.step(actions, step_time=30)
-IV-B UTracking Environment
+IV-B PyLrauv UTracking Environment
 UTrackingEnv is a Gym-like environment managing agent and target commands, communication, tracking, and RL tasks like observation building and reward calculations.
 Agents communicate via the Gazebo LRAUV Acoustic Comms Plugin and collect target ranges using the LRAUV Range Bearing Plugin.
-At the start of an episode, agents are spawned on the sea surface and targets at random depths, ensuring a minimum distance between each entity of 50 meters and maximum of 200.
+At the start of an episode, agents are spawned on the sea surface and targets at random depths, ensuring a minimum distance between each entity of 50 m and maximum of 200 m.
 Both agents and targets are LRAUV vehicles.
-Each step lasts 30 seconds in our experiments, at the beginning of which agents listen for range signals from targets, and at the end broadcast their locations and observations to other agents.
-Each agent maintain its own tracking model (LS or PF) for each target, updated with all the information received.
+Each step lasts 30 seconds in our experiments, at the beginning of which agents listen for range signals from targets, and at the end broadcast their locations and observations to other agents (communication phase).
+Each agent maintains its own tracking model (LS or PF) for each target, updated with all the information received.
 This ensures a genuine decentralization of the system.
-Targets are tracked in 2D space, as depth is typically known [ivanscience].
+Targets are tracked in 2D space, as depth is typically known [7].
 Observation Space
@@
 For other agents, these distances are calculated using the information they have received from them (if available), and for targets, the distances are calculated from the tracking predictions generated by their own tracking modules.
-[ivanscience] incorporates the range observations into observation vectors, which may be helpful, but may also cause agents to overfit to training dynamics and overlook tracking information.
+[7] incorporates the range observations into observation vectors, which may be helpful, but may also cause agents to overfit to training dynamics and overlook tracking information.
 To prevent this, we limit observations to tracking-derived values, ensuring agents rely on tracking data and its inherent errors.
@@
 $r_{i}^{follow}=\mathbb{I}\left\{d_{i}\leq\,d_{\min}\right\},$ (2)
-[…]
-$r=\frac{1}{N}\sum_{i=1}^{N}r_{i},$ (3)
-where $N$ is the number of targets to track, i.e.
+with the threshold set in our experiments to $50\,\text{m}$ (and $100\,\text{m}$ for fast targets).
+In both cases, the global reward to the system is given by $r=\frac{1}{N}\sum_{i=1}^{N}r_{i},$ where $N$ is the number of targets to track, i.e.
 the reward is normalized by the number of targets to be always in the interval $[0,1]$, so that learned value functions can be transferred across scenarios with different numbers of targets.
@@
 To simplify the action space of our multi-agent system, we keep the agents’ velocity constant and allow them to control only the rudder.
-The rudder is discretized into five values corresponding to equidistant angles ranging from $-0.24$ to $0.24$ radians.
+The rudder is discretized into five values corresponding to equidistant angles ranging from 0.24-0.24 to $0.24$ radians.
 Agents can adjust the rudder only to the next closest two angles, preventing abrupt changes in direction.
@@
 The episode ends when the maximum number of steps is reached.
-The environment is not truncated when crashes occur or targets are lost, as we experimentally observed that truncating episodes would create training instabilities.
-Why UTracking is a difficult environment?
-  • Partial observability: Agents detect targets only within 450 m via noisy acoustic signals and rely on intermittent and unstable communication (1500 m range) for shared tracking data.
-  • Stochasticity: Sensor noise, communication dropouts, environmental perturbations, and unpredictable target motions amplify uncertainty.
-  • Coordination challenge: Agents must coordinate almost perfectly especially when tracking multiple rapid-moving targets, since a lack of coordination could result in rapidly loosing a target.
-  • Long horizons: Hour-scale missions (corresponding to thousands of steps) require learning delayed consequences of actions, which can be difficult in Multi-Agent scenarios given their non-stationarity.
-Moreover, they need to be robust to error accumulation in tracking models.
-IV-C Environment Distillation: JaxLrauv UTracking
-LRAUV Sim simulates the dynamics of a LRAUV robot at the millisecond level.
-However, LRAUV vehicles take actions only every few seconds to minutes.
-Morover, the tracking missions we are interested in span several hours.
-For this reason, it would be beneficial to simulate the environment state at a much larger time scale.
-Formally, we are interested in modeling the new position in the 2D plane $\mathbf{p}_{t+1}\in\mathbb{R}^{2}$ after a delta time $\delta_{t}$ given the actual robot position $\mathbf{p}_{t}$, absolute speed $v$ and rudder angle $\gamma$.
-Instead of reinventing the wheel and re-building an accurate physical model which is already provided by Gazebo, we want to approximate the change in the robots’ heading $\delta_{\psi}$ in order to use a simplified trajectory model, i.e.
-$\mathbf{p}_{t+1}=\mathbf{p}_{t}+v\delta_{t}\begin{bmatrix}\cos(\psi_{t}+\delta_{\psi}),\sin(\psi_{t}+\delta_{\psi})\end{bmatrix}$.
-Fixing $v$ and $\delta_{t}$, we parametrize $\delta_{\gamma}$ in respect to the rudder angle $\theta(\gamma)$.
-To approximate $\theta(\gamma)$ we follow an iterative distillation process: we collect trajectories in PyLrauv, build a simplified model, train an agent on that, and collect new data using the trained agent to improve the model distillation.
-We found that using an ensemble of linear models it is possible to approximate $\theta(\gamma)$ with a mean absolute error below 0.015 radians and a global $R^{2}$ score of 0.99.
-Such a simple model fails to grasp long time-dependency, and diverges when used autoregressively to generate long trajectories.
-Nevertheless, it performs well for shorter trajectories and is extremely faster than, for example, a $LSTM$.
-To improve the robustness of the system, we introduce Gaussian noise with a standard deviation of 0.02 radians, forcing the trained agent to handle trajectory uncertainty.
+Why is UTracking a difficult environment?
+UTracking poses a challenging benchmark due to partial observability, where agents can only detect targets within 450 m using noisy acoustic signals and share data through intermittent communication limited to 1500 m.
+The environment is highly stochastic, with sensor noise, communication dropouts, environmental perturbations, and unpredictable target motions.
+Effective tracking requires near-perfect coordination among agents, especially when following multiple fast-moving targets, as poor coordination can quickly lead to target loss.
+Finally, the long horizons of hour-scale missions (thousands of steps) demand learning policies that handle delayed consequences and remain robust to error accumulation in non-stationary multi-agent settings.
+IV-C JaxLrauv UTracking
+LRAUV Sim simulates the dynamics of an LRAUV robot at the millisecond level.
+However, in practice, LRAUV vehicles take actions only every few seconds to minutes, while the tracking missions of interest span several hours.
+For this reason, it is beneficial to simulate the environment state at a much larger time scale.
+Formally, we are interested in modeling the new position in the 2D plane $\mathbf{p}_{t+1}\in\mathbb{R}^{2}$ after a time step $\delta_{t}$, given the current robot position $\mathbf{p}_{t}$, absolute speed $v$, and rudder angle $\gamma$.
+Instead of re-building an accurate physical model (already provided by Gazebo), we approximate the change in the robot’s heading, $\delta_{\psi}$, in order to use a simplified trajectory model: $\mathbf{p}_{t+1}=\mathbf{p}_{t}+v\delta_{t}\begin{bmatrix}\cos(\psi_{t}+\delta_{\psi}),\sin(\psi_{t}+\delta_{\psi})\end{bmatrix}$.
+Fixing $v$ and $\delta_{t}$, we parametrize $\delta_{\psi}$ with respect to the rudder angle through a function $\theta(\gamma)$.
+To approximate $\theta(\gamma)$, we collect trajectories in PyLrauv and train a supervised model that predicts $\delta_{\psi}$ from $\theta(\gamma)$.
+We found that using a linear model for each fixed combination of $v$ and $\delta_{t}$, $\theta(\gamma)$ can be approximated with a mean absolute error below 0.015 radians and a global $R^{2}$ score of 0.99.
+Based on this, we build an ensemble of linear models for different combinations of $v$ and $\delta_{t}$.
+While such a simple model cannot capture long-term dependencies and diverges when used autoregressively for long trajectories, it performs well for shorter trajectories and is much faster than, for example, an LSTM or a handwritten physical system.
+To improve robustness, we introduce Gaussian noise with a standard deviation of 0.02 radians, forcing the trained agent to handle trajectory uncertainty.
 In addition, we equip the agents with recurrent mechanisms.
-We found that training in this way in the simplified environment, the learned policies are directly transferable to Gazebo.
-The rest of the environment, including measurement noise, the probability of communication loss, partial observability and collision handling, is implemented manually following the observed statistics from the collected data.
-The final simplified environment is implemented in JAX to enable an end-to-end GPU-based RL pipeline, and is called JaxLrauv UTracking.
-The speedups achieved in different configurations of agents and targets are presented in LABEL:tab:speedup.
-These speedups allow to train a single-target tracking model that can be directly deployed in LRAUV Sim in as little as 10 minutes.
-TABLE I: Speedups Obtained with the Distilled Environment.
+We found that training in this simplified environment leads to policies that transfer directly to Gazebo.
+The rest of the environment, including measurement noise, communication loss, partial observability, and collision handling, is implemented manually based on data collected in Gazebo.
+The final simplified environment is implemented in JAX, enabling an end-to-end GPU-based RL pipeline.
+To further improve training efficiency, we implement a Particle Filter in JAX, allowing the tracking phase to also benefit from GPU acceleration.
+By nesting vectorized functions, we update all particles across targets, agents, and parallel environments in a single pass, achieving both high speed and precise tracking during training.
+The same vectorized PF algorithm is also integrated into PyLrauv for consistency.
+The speedups achieved across different agent–target configurations are reported in LABEL:tab:speedup.
+These gains make it possible to train a single-target tracking model that can be directly deployed in LRAUV Sim in as little as 10 minutes.
+TABLE I: Speedups Obtained with the Accelerated Environment. Steps Per Second (SPS) and relative speedup.
 Config Pylrauv JaxLrauv 1 Envs JaxLrauv 128 Envs JaxLrauv 1024 Envs
@@
 5A,5T 0.2 936 4751 3469 17607 3574 18142
-IV-D Vectorized Particle Filter
-While PF represent a state-of-the-art method for underwater tracking, their computational expense renders them inefficient for training, particularly in multi-agent scenarios.
-To mitigate this, [ivanscience] suggests employing Least Squares (LS) during training and reserving PF for testing.
-However, LS is only stable for static or slow-moving targets, limiting its scalability for faster targets and potentially intensifying the inherent non-stationarity of MARL.
-To facilitate efficient training, we implement a PF in JAX, utilizing vmap for GPU parallelization.
-By nesting vmapped functions, during training we update all particles across targets, agents, and parallel environments in a single pass, reducing the speed loss to only 30% compared to LS.
-As demonstrated in Figure 3, training with LS fails for targets moving faster than one-third of the agent’s velocity, whereas PF maintains stability.
-Figure 3: Using Particle Filter or Least Squares to train an agent to follow a target at different velocities.
-IV-E TransfMAPPO
-The complexity of training multi-agent system with RL typically scales with the number of agents involved.
-To address this, we introduce TransfMAPPO, inspired by TransfQMix [transfqmix], which maintains the core concepts of the latter but uses PPO as the RL backbone instead of DQN, since PPO is more natural for vectorized environments.
-In TransfMAPPO, we implement a Transformer Actor and Critic that process the features describing the observed entities as vertices of a coordination latent graph learned through self-attention.
-The resulting policy and Value function are therefore invariant from the number of entities observed, allowing curriculum learning procedure described in subsection IV-F.
-More formally, the Transformer Critic observes the state of the environment represented as $n$ entities each one described by $z$ features, which in our case are the position, velocity and angles of the agents and targets.
-The Transformer Agent also observes the $n$ entities, but from its own partial observability perspective, which is constraint by the acoustic communication protocols defined in section IV.
-An overview of TransfMAPPO can be seen in Figure 4.
-Following [gallici2024simplifyingdeeptemporaldifference], we found that normalizing the entire network with LayerNormalization greatly improves stability and performances.
+IV-D TransfMAPPO
+The complexity of training multi-agent systems with RL typically scales with the number of agents involved.
+To address this, we introduce TransfMAPPO, inspired by TransfQMix [15], which preserves the core concepts of the latter but uses PPO instead of DQN as the RL backbone, since PPO is naturally suited for vectorized environments.
+To leverage Transformers, we formalize the problem as learning a latent coordination graph through self-attention, where only the vertices of the graph are known.
+Specifically, the $n$ vertices (or tokens) correspond to the agents and targets present in the environment, each represented by $z$ features such as position, velocity, and orientation.
+This representation is permutation-invariant, allowing the Transformer to learn a coordination graph that is also invariant to the number of agents and targets—much like an LLM can learn semantic attention patterns over sentences of varying lengths.
+We distinguish between two types of latent representations: (1) the local latent graphs which can be learned by each agent aggregating noisy and potentially incomplete information from onboard sensors and inter-agent communications, and (2) the global latent graph learnable using the absolute states of all agents and targets, which is available only during training in simulation.
+Accordingly, we use two Transformer networks Figure 3: the TransformerAgents, deployed in a decentralized fashion (one per vehicle), which learns the local coordination graph to control the agents using only their local observations; and the TransformerCritic, which learns a value function for the joint action–state pair using the global information, and is employed for centralized training.
+The use of Transformers in both actor and critic improves coordination representation and yields policies and value functions that are invariant to the number of entities.
+This property enables the curriculum learning procedure described in subsection IV-E.
+Following [31], we found that normalizing the entire network with LayerNormalization greatly improves stability and performances.
 Figure 3: Overview of TransfMAPPO architecture.
-IV-F Curriculum Learning
-Taking advantage of our Transformer-based architecture, we employ curriculum learning to progressively learn Multi-Agent general policies starting from a Single-Agent policy (see Figure 5).
+IV-E Curriculum Learning
+Taking advantage of our Transformer-based architecture, we employ curriculum learning to progressively learn Multi-Agent policies starting from a Single-Agent policy (see Figure 4).
 We begin with a large pre-training phase in which we train a single agent to maximize the tracking reward described in subsection IV-A with respect to an erratic-moving fast target (0.6x of its speed).
-We train for $10^{9}$ timesteps, corresponding to 1.5 days of compute (that would take months in Gazebo).
+We train for $1^{10}$ timesteps, corresponding to 1.5 days of compute (that would take months in Gazebo).
 In this phase, we maintain a short horizon (128 steps) before resetting the environment, as it significantly aids learning.
-The resulting policy is able to track a fast target with an error below 35 meters, but it fails to generalize to longer horizons.
-We therefore fine-tune the policy in progressively longer episodes (256, 512, and 1024 steps) for $10^{7}$ timesteps, until we observe that the policy can perfectly track a fast target for more than 10k steps, corresponding to 3.4 days of real-time tracking.
+The resulting policy is able to track a fast target with an error below 35m, but it fails to generalize to longer horizons.
+We therefore fine-tune the policy in progressively longer episodes (256, 512, and 1024 steps) for $1^{8}$ timesteps, until we observe that the policy can perfectly track a fast target for more than 10k steps, corresponding to more than 3 days of real-time tracking.
 From this horizon-invariant base model, we transition to a multi-agent phase, where we fine-tune the policy for cooperative tasks.
@@
 Given that these two scenarios pose different cooperation and credit assignment problems, we reset the Critic parameters and maintain separate sets of fine-tuned agents for the two multi-agent fine-tuning branches, resulting in a final multi-target follow model and a multi-robot tracking model.
-In both cases, we fine-tune the model to cooperate with progressively more agents (5 and 3 at maximum, respectively) for a maximum $20^{8}$ timesteps.
-We observe that both resulting models can cooperate with a variable number of agents.
-In this phase, we also note that maintaining long environment horizons negatively affects training, so we reduce the horizon back to 256 steps.
-Unfortunately, while the resulting multi-agent tracking model retains the horizon invariance of the pre-trained single-agent model, the multi-target follow model loses some invariance in comparison.
+In both cases, we fine-tune the model to cooperate with progressively more agents (5 and 3 at maximum, respectively) for a maximum $2^{9}$ timesteps.
 Figure 4: Overview of our Curriculum Learning Procedure.
@@
 V-A Multi-Robot Tracking
-The effectiveness of our curriculum method is demonstrated in Figure 6, where we compare the training curve of fine-tuning the model for the task of tracking a very fast target (0.8x the agents’ velocity) with 3 agents against training from scratch on the same task.
-The superior performance of curriculum learning is also evident in the reduced tracking error.
-In the same plot, we also include the training curve of single-agent PPO, which achieves less than half the returns of the multi-agent system.
+In Figure 5, we show the training curve for the most complex multi-robot tracking task we consider, i.e., three agents tracking a very fast target (0.8× the agents’ velocity).
+Specifically, we present results for training standard MAPPO (with an RNN) and TransfMAPPO from scratch, as well as the result of fine-tuning TransfMAPPO on this task following the curriculum procedure described in subsection IV-E.
+Note that training Transformers from scratch on this task is not more effective than using an RNN; however, the benefits of curriculum learning are evident in terms of expected returns and reduced tracking error.
 Figure 5: Training multiple agents to track a very fast target.
-Figure 7: Multi-Robot Tracking Evaluation. The highglighted area represents the values seen in training.
-To validate our final multi-robot tracking model based on TransfMAPPO, we conduct a comprehensive validation analysis using JaxLrauv Figure 10.
-Every data point in the validation phase is averaged across 1000 episodes, each with a length of 256 steps, following a target moving at 0.5x the agents’ velocity (unless otherwise specified).
-Specifically, we test how the tracking error is affected under three conditions: (1) when the target moves faster, (2) when its motion becomes more unpredictable (defined by the frequency of direction changes), and (3) when the agents’ motion is perturbed (simulating external forces such as ocean currents).
-In the last sub-figure we illustrate how the multi-agent system’s performance scales naturally with the number of agents used to track the target.
-Notably, even using only 2 agents significantly enhances the system’s robustness compared to a single-agent approach.
-We also highlight that the tracking error decreases monotonically with the number of agents, with 3 agents representing the sweet spot, achieving negligible collision probability while maintaining optimal tracking performance.
+Figure 6: Multi-Robot Tracking Evaluation. The highlighted area represents the values seen in training.
+In Figure 6, we perform a thorough evaluation of the final multi-target follow model obtained through curriculum training.
+Each data point in the validation phase is averaged over 1000 episodes, each lasting 256 steps, with the target moving at 0.5× the agents’ velocity (unless otherwise specified).
+Specifically, we assess how the tracking error is affected under three conditions: (1) when the target moves faster, (2) when its motion becomes more unpredictable (defined by the frequency of direction changes), and (3) when the agents’ motion is perturbed (simulating external forces such as ocean currents).
+In the last sub-figure, we illustrate how the performance of the multi-agent system scales with the number of agents used to track the target.
+As expected, even using only two agents significantly improves the system’s robustness compared to a single-agent approach.
+Robustness increases proportionally with the number of agents, both in terms of resistance to variations in target velocity, target motion (i.e., the interval of target direction changes), and perturbation noise (additional noise added to agents’ trajectories).
+We also highlight that the tracking error decreases monotonically with the number of agents, with three agents representing a sweet spot: achieving negligible collision probability while maintaining optimal tracking performance.
 V-B Multi-Target Tracking
@@
 Figure 7: Training Multiple Agents to Follow Fast Targets.
-To further test the effectiveness of TransfMAPPO, we train it from scratch alongside MAPPO in a simpler configuration where targets move slower (0.3x the agents’s speed) and in a more predictable manner.
+To further test the effectiveness of TransfMAPPO, we train it from scratch alongside MAPPO in a simpler configuration where targets move slower (0.3x the agents’s speed) and in a more predictable manner (Figure 8).
 In this case, TransfMAPPO and MAPPO achieve similar scores with 2 agents, but MAPPO’s performance drops significantly with 3 agents.
-Surprisingly, TransfMAPPO learns very effectively even from scratch in the 5-agent-5-target scenario, confirming the sample efficiency obtained by decoupling a cooperation policy using Transformers.
+Surprisingly, TransfMAPPO learns very effectively even from scratch in the 5-agent-5-target scenario, confirming the sample efficiency obtained by learning a coordination graph using Transformers.
 Figure 8: Training Multiple Agents to Follow Slower Targets.
 V-C Results in Gazebo
+TABLE II: Evaluation in JaxLrauv and PyLrauv of the curriculum training policies (* except for first line which refers to 10 minutes training).
+Configuration Avg. Agent-Target Distance Avg. Tracking Error Probability of Collision (%) Probability of Loosing a Target (%)
+JaxLrauv PyLrauv JaxLrauv PyLrauv JaxLrauv PyLrauv JaxLrauv PyLrauv
+1 Agent, 1 Target (Slow)* 55.20 ± 30.11 59.10 ± 32.29 5.12 ± 3.1 5.89 ± 3.7 0.00 0.00 0.00 0.00
+1 Agent, 1 Target (Fast) 110.23 ± 58.90 121.01 ± 62.11 17.40 ± 21.5 20.33 ± 28.38 0.00 0.00 5.00 7.14
+3 Agents, 1 Target (Very Fast) 200.11 ± 85.70 217.64 ± 89.85 2.65 ± 5.30 3.03 ± 6.17 0.0 15.32 0.00 0.00
+3 Agents, 3 Targets (Moderate) 42.15 ± 25.00 45.30 ± 26.11 4.85 ± 4.20 5.29 ± 4.88 0.5 5.00 3.50 5.00
+5 Agents, 5 Targets (Moderate) 46.12 ± 28.75 49.24 ± 30.92 3.80 ± 4.00 4.25 ± 4.57 2.1 10.00 4.10 5.26
 We conducted a series of final experiments in Gazebo to evaluate the performance of our trained agents in a high-fidelity simulation environment.
@@
   1. 1.
-    Tracking a slow target (0.3$\times$ the agent’s velocity) for 300 steps using a single robot trained for 10 minutes in JaxLrauv.
+[…]
   2. 2.
-    Tracking a fast target (0.6$\times$ the agent’s velocity) for 1000 steps (equivalent to over 8 hours of real-time tracking) using our horizon-invariant base model.
+[…]
   3. 3.
-    Tracking a very fast target (0.8$\times$ the agent’s velocity) for 1000 steps using a multi-robot tracking fine-tuned model with 3 agents.
+[…]
   4. 4.
-    Tracking 3 targets simultaneously moving at 0.5$\times$ the agents’ velocity using a multi-target model for 300 steps.
+[…]
   5. 5.
-    Tracking 5 targets simultaneously moving at 0.5$\times$ the agents’ velocity using a multi-target model for 300 steps.
+[…]
 For the multi-target experiments, we reduced the number of steps to 300 due to the computational burden of simulating multiple robots in Gazebo.
 The results of these experiments are presented in Table II.
+For completeness, we also include in this table the results obtained in JaxLrauv with the same configurations, so that it is possible to appreciate how the learned policy behaves in a very similar way in both simulators, resulting in comparable evaluation measures.
 Our findings demonstrate that even with only 10 minutes of training, we were able to develop an effective model capable of maintaining target tracking without losing the target.
-The multi-robot tracking model achieved a promising average tracking error of 3 meters, even when following a very fast target.
+The multi-robot tracking model achieved a promising average tracking error of 3m, even when following a very fast target.
 This performance highlights a clear advantage over single-agent trackers following slower targets.
@@
 This underscores a critical area for future work to enhance the systems’ safety before deploying them in real-world marine missions.
-TABLE II: Experimental Results in Gazebo
-Config Avg. Agent-Target Distance Av. Tracking Error Prob. Collision (%) Prob. Lost (%)
-1A,1T (Slow, 10Mins Train) 59.10 ± 32.29 5.89 ± 3.7 0.00 0.00
-1A,1T (Fast) 121.01 ± 62.11 20.33 ± 28.38 0.00 7.14
-3A,1T (Very Fast) 217.64 ± 89.85 3.03 ± 6.17 15.32 0.00
-3A,3T (Moderate) 45.30 ± 26.11 5.29 ± 4.88 5.00 5.00
-5A,5T (Moderate) 49.24 ± 30.92 4.25 ± 4.57 10.00 5.26
+Finally, in Figures 9 to 12 we present some of the trajectories obtained during the final evaluation, together with zoomed-in views of the emergent coordination strategies learned.
 VI Conclusions
 We presented a method for scaling up Multi-Agent Reinforcement Learning (MARL) in the context of underwater tracking, demonstrating how it is possible to efficiently train MARL models purely on GPUs and deploy them in high-fidelity simulators, where MARL training would otherwise be extremely expensive or infeasible.
-Although this method relied on manual adjustments during the environment distillation phase, our overall pipeline represents a promising direction for closing the sim-to-real gap in MARL.
+Our overall pipeline represents a promising approach for closing the sim-to-real gap in MARL.
 A natural next step for this work is to enhance the safety of the multi-agent system and test it in real-world sea missions.
-Figure 10: Multi-agent episodes collected in Gazebo.
-(a): Single-agent tracking, target velocity 0.66× of agent velocity.
-Video.
-(b): Multi-agent tracking, target velocity 0.88x of agents velocity.
-Video.
-Figure 11: Agents trajectory according to target speed. Notice the change in the curvature of the agent’s trajectory to optimally track targets moving at different speeds.
+Figure 9: (a): Single-agent tracking, target velocity 0.66× of agent velocity. Video. (b): Multi-agent tracking, target velocity 0.88x of agents’ velocity. Video.
+Figure 10: Agents’ trajectory according to target speed. Notice the change in the curvature of the agent’s trajectory to optimally track targets moving at different speeds.
 Figure 11: Coordination when tracking a very fast target.
-Notice how agents react to a rapid change of direction of the target by waiting each other in order to not loose communication and then rapidly moving together.
-Figure 13: Coordination in multi-target setting.
-Notice how agents resolve the ”traffic” problem by making circles to wait for other agents’ passing, and one target is dynamically reassigned from agent 3 to agent 1.
+Notice how agents react to a rapid change of direction of the target by waiting for each other in order not to lose communication and then rapidly moving together.
+Figure 12: Coordination in a multi-target setting.
+Notice how blue and orange agents resolve the ”traffic” problem by making circles to wait for other agents’ passing, and one target is dynamically reassigned from one agent to another.
 ACKNOWLEDGMENT
@@
 G.
-was partially founded by the FPI-UPC Santander Scholarship FPI-UPC_93.
-This work also acknowledges the ’Severo Ochoa Centre of Excellence’ accreditation (CEX2019-000928-S) and the Barcelona Super Computer Center.
-[…]
+was partially funded by the FPI-UPC Santander Scholarship FPI-UPC_93.
+This work also acknowledges the ‘Severo Ochoa Centre of Excellence’ accreditation (CEX2024-001494-S) from AEI 10.13039/501100011033 and the Barcelona Supercomputer Center.